OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. -3, -15, -75, -375, ...

OpenStudy (anonymous):

I got an = -3 • 5n - 1

OpenStudy (anonymous):

the problem (assuming that -3 is referred to as n= 0) is that you don't get the first term in the sequence. one thing to try here it to use some number to a power (because any number to the 0th power is one)

OpenStudy (anonymous):

Continue...wait, is my answer incorrect?

OpenStudy (anonymous):

it is incorrect, for n = 0, you got -3( 5 * 0) - 1 0 - 1 if you put in n = 1, you get -3 (5 * 1) -1 -15 - 1

OpenStudy (anonymous):

oh wait, I misread the sequence. you go tit right

OpenStudy (anonymous):

it is a G.P with C.R=-15/-3=5 tn =ar^(n-1)

OpenStudy (anonymous):

a is first term, r is a C.R

OpenStudy (anonymous):

tn=-3(5)^(n-1)

OpenStudy (anonymous):

Thank You, @alexandercpark and @surjithayer :)